What is the molar concentration of aki’s disinfectant if she dissolved 5Tbsp. Of NaCIO (sodium hypochlorite)in 3.8 liters of H2o(water)


1
Expert's answer
2021-11-24T12:34:05-0500

5 tablespoons = 73.9ml

73.9ml = 73.9 cm³ = 73.9 × 10-3dm³


Since, 22.4 dm³ = 1 mole

"\therefore" 73.9 × 10-3 dm³ = 0.0033 moles


molar concentration = number of moles/volume

= 0.0033/3.8 = 0.0087M"."